OpenStudy (anonymous):

use the basic identities to simplify the expression cos^2theta -(csc theta-1)(csc theta +1)

OpenStudy (abb0t):

Csc = 1/sin Cos(2x) = cos^2(x) - sin^2(x) = 1/2 [1+cos(x)]

OpenStudy (anonymous):

thank you!

OpenStudy (preetha):

@minabina678 Please give @abb0t a medal for his help!

OpenStudy (abb0t):

Aww! Thanks @Preetha :)

OpenStudy (dumbcow):

how far do they want it simplified? do you have the answer? here is what i see \[=\cos^{2} \theta - (\csc^{2} \theta -1)\] \[=\cos^{2} \theta - \cot^{2} \theta\] \[=-\cos^{2} \theta \cot^{2} \theta\]